### v1

@@ 2 Preliminaries
In this section, we discuss the tools needed for our proposed methods in Section 3. Many of the methods described here are in the randomized numerical linear algebra literature. For an in-depth review, we refer to [28, 44, 58]. First, in Section 2.1, we discuss the core ingredient in many randomized numerical linear algebra algorithms, which are random embeddings. Following this, in Section 2.2, we review pivoting on a random sketch, which efficiently computes the indices for the CUR decomposition, and in Section 2.3, we review a fast randomized algorithm for computing the numerical rank of a matrix. Lastly, in Section 2.4, we discuss an efficient method for computing the Frobenius norm of a matrix, which is used for error estimation for AdaCUR in Section 3.
@@ 2.1 Random embeddings
for every $x\in\mathbb{R}^{n}$. Therefore $\Gamma$ is a linear map, which preserves the $2$-norm of every vector in a given subspace. A random embedding is a subspace embedding drawn at random that satisfies (4) for all $A$ with high probability. A typical application of random embeddings is in matrix sketching, a technique for dimensionality reduction. Matrix sketching compresses the original matrix into a smaller-sized matrix while retaining much of its original information. For example, for a matrix $A\in\mathbb{R}^{m\times n}$ of rank $r\ll\min\{m,n\}$, $\Gamma A\in\mathbb{R}^{s\times n}$ where $r\leq s\ll\min\{m,n\}$ is a (row) sketch of the matrix $A$. Here, the integer $s$ is called the sketch size.
There are a few important examples of random embeddings such as Gaussian embeddings [28, 44], subsampled randomized trigonometric transforms (SRTTs) [5, 54] and sparse embeddings [12, 13, 47]. We focus on Gaussian embeddings in this work. A Gaussian embedding $\Gamma\in\mathbb{R}^{s\times m}$ has i.i.d. entries $\Gamma_{ij}\sim\mathcal{N}(0,1/s)$. Here, $s=\Omega(r/\epsilon^{2})$ for theoretical guarantees, but $s=r+\Omega(1)$ works well in practice. The cost of applying a Gaussian embedding to a matrix $A\in\mathbb{R}^{m\times n}$ is $\mathcal{O}(sT_{A})$ where $T_{A}$ is the cost of matrix-vector multiply with $A$ or $A^{T}$. Other random embeddings provide a similar theoretical guarantee but generally require a larger sketch size, say $s=\Omega(r\log r/\epsilon^{2})$. However, they are usually cheaper to apply, for example, SRTTs cost $\mathcal{O}(mn\log s)$ [59].
@@ 2.2 Pivoting on a random sketch
Pivoting on a random sketch [19, 22, 57] is an attractive and efficient method for finding a spanning set of columns and/or rows of a matrix. This will be the core part for computing the CUR decomposition. This method has two main steps: sketch and pivot. Let $A\in\mathbb{R}^{m\times n}$ be a matrix and $r$ be the target rank, then the basic idea works as follows for column indices:
@@ 2.3 Randomized rank estimation
Randomized rank estimation [2, 45] is an efficient tool for finding the $\epsilon$-rank of a matrix $A\in\mathbb{R}^{m\times n}$. As we often require the target rank as input, for example in Algorithm 1, we would like an efficient method for computing the $\epsilon$-rank of a matrix. This method also relies on sketching, but unlike pivoting on a random sketch, it involves sketching $\Gamma_{1},\Gamma_{2}$ from both sides of the matrix. The idea is to approximate the $\epsilon$-rank of $A$ using
@@
where $\operatorname{rank}_{\epsilon}(B)$ is the $\epsilon$-rank of the matrix $B$, i.e. the number of singular values larger than $\epsilon$, and $\Gamma_{1}A\Gamma_{2}\in\mathbb{R}^{s\times 2s}$ with $s\gtrsim\operatorname{rank}_{\epsilon}(A)$. Here, $s\gtrsim\operatorname{rank}_{\epsilon}(A)$ is achieved by gradually increasing $s$ by appending to the sketches until $\sigma_{\min}(\Gamma_{1}A\Gamma_{2})<\epsilon$; see [45] for details. The overall cost of the algorithm is $\mathcal{O}(sT_{A}+ns\log s+s^{3})$ where $\mathcal{O}(sT_{A})$ is the cost of forming the Gaussian sketch, $\mathcal{O}(ns\log s)$ is the cost of forming the SRTT sketch and $\mathcal{O}(s^{3})$ is the cost of computing the singular values of $\Gamma_{1}A\Gamma_{2}$.
@@ 2.4 Randomized norm estimation
Randomized norm estimation [26, 44] is an efficient method for finding an approximation to a norm of a matrix. In this work, we focus on the Frobenius norm as it is a natural choice for low-rank approximation and easier to estimate. In order to monitor how well our algorithm performs, we need an efficient way to estimate the error. Let $A\in\mathbb{R}^{m\times n}$ be a matrix and $\hat{A}_{k}\in\mathbb{R}^{m\times n}$ be a rank-$k$ approximation to $A$. Then we would like to estimate $\left\lVert A-\hat{A}_{k}\right\rVert_{F}$. As before, we can sketch to get an approximation to $A-\hat{A}_{k}$ and then compute the Frobenius norm. It turns out that the sketch size of $O(1)$, say $5$, suffices. More specifically, we have the following theorem from [26].
@@
The small sketch can be used to obtain an additional set of row and column indices by using pivoting on it, similarly to Algorithm 1. By adding these additional indices, we can improve the accuracy of the CUR decomposition. This observation will be used later in AdaCUR.
@@ 3 Proposed method
Let us first state the problem. Let $A(t)\in\mathbb{R}^{m\times n}$ be a parameter-dependent matrix with $t\in D$ where $D\subseteq\mathbb{R}^{d}$ is a compact domain and $t_{1},t_{2},...,t_{q}\in D$ be a finite number of distinct sample points ordered in some way, e.g., $t_{1}<t_{2}<\cdots<t_{q}$ for $t\in\mathbb{R}$. Then find a low-rank approximation of $A(t_{1}),A(t_{2}),...,A(t_{q})$. [^7] Although our algorithms are presented for a finite number of sample points, they are likely to work on a continuum, e.g. by choosing the indices for the closest point $t_{i}$ for each $t\in[t_{1},t_{q}]$.
@@
The goal of this work is to devise an efficient algorithm that reuse the indices to their fullest extent in the CUR decomposition as we iterate through the parameter values. To see how we can use the same indices for different parameter values, recall (2) with oversampling from [50]:
(9) $\left\lVert A-A_{I\cup I_{0},J}\right\rVert_{F}\leq\left\lVert Q_{C}(I\cup I_{0},:)^{\dagger}\right\rVert_{2}\left\lVert Q_{X}(J,:)^{-1}\right\rVert_{2}\left\lVert A-AX^{\dagger}X\right\rVert_{F}.$
@@
Here, $I_{0}$ is an extra set of row indices resulting from oversampling that is distinct from $I$.
The bound (9) leads us to two observations. First, when we have two matrices, say $A(t_{1})$ and $A(t_{2})$ with $r$-dominant right singular vectors $V_{1}\in\mathbb{R}^{n\times r}$ and $V_{2}\in\mathbb{R}^{n\times r}$, respectively, we get
(10) $\left\lVert A(t_{1})-A(t_{1})_{I_{*}J}\right\rVert_{F}\leq\left\lVert Q_{C_{1}}(I_{*},:)^{\dagger}\right\rVert_{2}\left\lVert V_{1}(J,:)^{-1}\right\rVert_{2}\left\lVert A(t_{1})-\llbracket{A(t_{1})}\rrbracket_{r}\right\rVert_{F}$
and
(11) $\left\lVert A(t_{2})-A(t_{2})_{I_{*}J}\right\rVert_{F}\leq\left\lVert Q_{C_{2}}(I_{*},:)^{\dagger}\right\rVert_{2}\left\lVert V_{2}(J,:)^{-1}\right\rVert_{2}\left\lVert A(t_{2})-\llbracket{A(t_{2})}\rrbracket_{r}\right\rVert_{F}$
where $I_{*}=I\cup I_{0}$, $C_{1}=A(t_{1})(:,J)$ and $C_{2}=A(t_{2})(:,J)$. Therefore when the first two terms of the right-hand sides of (10) and (11) are modest, the same indices $I_{*}$ and $J$ can be used for the CUR decomposition of $A(t_{1})$ and $A(t_{2})$. This happens when the set of column indices $J$ captures the important rows of both $V_{1}$ and $V_{2}$ and the set of row indices $I$ captures the important rows of the orthonormal matrices spanning $C_{1}$ and $C_{2}$. Interestingly, this situation occurs quite frequently in applications even when $\left\lVert A(t_{1})-A(t_{2})\right\rVert_{2}$ is not small; see Section 4. This observation motivates us to check if the previously obtained indices can still be used before recalculating them entirely.
@@ 3.1 Computing indices from scratch
In the next two sections, we discuss our algorithms, AdaCUR and FastAdaCUR. For both algorithms, we need to start with an index set for the CUR decomposition. This can sometimes be done offline, for example, if we are solving a Matrix PDE, we can compute the initial set of indices for the CUR decomposition from the initial condition matrix. In this work, we use Algorithm 2 to get a set of indices from scratch. The procedure is to first approximate the $(\epsilon/\sqrt{n})$-rank of the initial matrix $A(t_{1})$ using randomized rank estimation and then apply pivoting on a random sketch with the estimated $(\epsilon/\sqrt{n})$-rank to get the initial set of indices. Here, we approximate the $(\epsilon/\sqrt{n})$-rank [^8] The tolerance parameter for randomized rank estimation can be set to slightly less than $\epsilon/\sqrt{n}$, such as $0.5\epsilon/\sqrt{n}$, to accommodate for both the randomization error and the quasi-optimal error of the CUR decomposition. to ensure a relative error of $\epsilon$ for the CUR decomposition in the Frobenius norm. The procedure for computing indices from scratch will be used for getting an initial set of indices for AdaCUR and FastAdaCUR, but also in AdaCUR when the error becomes too large that we need to recompute the indices altogether from scratch.
@@ 3.2 AdaCUR algorithm for accuracy
AdaCUR is an algorithm aimed at efficiently computing an accurate low-rank approximation to a given tolerance for a parameter-dependent matrix. An overview of AdaCUR goes as follows. The algorithm starts by computing the initial set of indices for $A(t_{1})$ as discussed in Section 3.1. For subsequent parameter values, i.e., $A(t_{j})$ for $j=2,...,q$, we first verify whether the indices obtained from the previous parameter value are able to meet some given error tolerance, using an estimate of $\|A-CU^{\dagger}R\|_{F}$. If so, we continue to the next parameter value. Should the indices fail to meet the tolerance, we make low-cost minor modifications to the indices and test whether the adjustments satisfy the error tolerance. If this is still unsuccessful, we recompute the indices entirely from scratch and move onto the next parameter value. Minor modifications include adding a small fixed number of indices, reordering them based on their importance and removing some if necessary. The details are in the following paragraph. The addition and removal of indices, as well as recomputation of indices using rank estimation if necessary, make the algorithm rank-adaptive and computing the relative error at each instance make the algorithm certifiable. AdaCUR is presented in Algorithm 5.
@@
12:        $[I_{1},J_{1}]=\mathtt{Rand\_Pivot}(A(t_{j})([m]-I,:),E_{s},s)$triangleright Algorithm 1 on unchosen indices
@@
14:        $[\sim,\sim,I_{2}]=\mathtt{sRRQR}\left(A(t_{j})\left(I\cup I_{0},J\right)^{T}\right)$ 999$[Q,R,I]=\mathtt{sRRQR}(A)$ performs strong rank-revealing QR factorization [27] on $A$ where $Q$ is an orthonormal matrix, $R$ is an upper triangular matrix and $I$ is a set of pivots.
@@
AdaCUR (Algorithm 5) takes the parameter-dependent matrix $A(t)$, evaluation points $t_{1},...,t_{q}\in D$, error tolerance $\epsilon$ and error sample size $s$, oversampling parameter $p$ as input. The algorithm produces the low-rank factors $C(t_{j}),R(t_{j}),U(t_{j})$ as output such that the factors are a subset of columns, rows and their intersection of $A(t_{j})$, respectively, and $A(t_{j})\approx C(t_{j})U(t_{j})^{\dagger}R(t_{j})$. If the entries of the parameter-dependent matrix can be computed quickly afterwards, the algorithm can be modified to output only row and column indices, saving storage. AdaCUR is quite involved, so we break it down into two core parts: (i) error estimation using previous indices (lines 55558888) and (ii) low-cost minor modifications (lines 1010101019191919). The other lines involve computing the indices from scratch, which follow the same discussion as Section 3.1.
(i) Initial error estimation

### v2

@@ 2 Preliminaries
In this section, we discuss the tools needed for our proposed methods introduced in Section 3. Many of the methods described here are in the randomized numerical linear algebra literature. For an in-depth review, we refer to [28, 44, 58]. First, in Section 2.1, we discuss the core ingredient in many randomized numerical linear algebra algorithms, which are random embeddings. Following this, in Section 2.2, we review pivoting on a random sketch, which efficiently computes the indices for the CUR decomposition, and in Section 2.3, we review a fast randomized algorithm for computing the numerical rank of a matrix. Lastly, in Section 2.4, we discuss an efficient method for computing the Frobenius norm of a matrix, which is used for error estimation for AdaCUR in Section 3.
@@ 2.1 Random embeddings
for every $x\in\mathbb{R}^{n}$. Therefore, $\Gamma$ is a linear map which preserves the $2$-norm of every vector in a given subspace. A random embedding is a subspace embedding drawn at random that satisfies (7) for all $A$ with high probability. A typical application of random embeddings is in matrix sketching, a technique for dimensionality reduction. Matrix sketching compresses the original matrix into a smaller-sized matrix while retaining much of its original information. For example, for a matrix $A\in\mathbb{R}^{m\times n}$ of rank $r\ll\min\{m,n\}$, $\Gamma A\in\mathbb{R}^{s\times n}$ where $r\leq s\ll\min\{m,n\}$ is a (row) sketch of the matrix $A$. Here, the integer $s$ is called the sketch size.
There are a few important examples of random embeddings such as Gaussian embeddings [28, 44], subsampled randomized trigonometric transforms (SRTTs) [5, 54] and sparse embeddings [12, 13, 47]. We focus on Gaussian embeddings in this work. A Gaussian embedding $\Gamma\in\mathbb{R}^{s\times m}$ has i.i.d. entries $\Gamma_{ij}\sim\mathcal{N}(0,1/s)$. Here, $s=\Omega(r/\epsilon^{2})$ for theoretical guarantees, but $s=r+\Omega(1)$ works well in practice. The cost of applying a Gaussian embedding to a matrix $A\in\mathbb{R}^{m\times n}$ is $\mathcal{O}(sT_{A})$ where $T_{A}$ is the cost of matrix-vector multiply with $A$ or $A^{T}$. Other random embeddings provide a similar theoretical guarantee but generally require a larger sketch size, say $s=\Omega(r\log r/\epsilon^{2})$. However, they are usually cheaper to apply; for example, SRTTs cost $\mathcal{O}(mn\log s)$ [59].
@@ 2.2 Pivoting on a random sketch
Pivoting on a random sketch [19, 22, 57] is an attractive and efficient method for finding a spanning set of columns and/or rows of a matrix. This will be the core part for computing the CUR decomposition. This method has two main steps: sketch and pivot. Let $A\in\mathbb{R}^{m\times n}$ be a matrix and $r$ be the target rank. Then the basic idea works as follows for column indices:
@@ 2.3 Randomized rank estimation
Randomized rank estimation [2, 45] is an efficient tool for finding the $\epsilon$-rank of a matrix $A\in\mathbb{R}^{m\times n}$. As we often require the target rank as input, for example in Algorithm 1, we would like an efficient method for computing the $\epsilon$-rank of a matrix. This method also relies on sketching but, unlike pivoting on a random sketch, it involves sketching $\Gamma_{1},\Gamma_{2}$ from both sides of the matrix. The idea is to approximate the $\epsilon$-rank of $A$ using
@@
where $\operatorname{rank}_{\epsilon}(B)$ is the $\epsilon$-rank of the matrix $B$ (i.e., the number of singular values larger than $\epsilon$) and $\Gamma_{1}A\Gamma_{2}\in\mathbb{R}^{s\times 2s}$ with $s\gtrsim\operatorname{rank}_{\epsilon}(A)$. Here, $s\gtrsim\operatorname{rank}_{\epsilon}(A)$ is achieved by gradually increasing $s$ by appending to the sketches until $\sigma_{\min}(\Gamma_{1}A\Gamma_{2})<\epsilon$; see [45] for details. The overall cost of the algorithm is $\mathcal{O}(sT_{A}+ns\log s+s^{3})$ where $\mathcal{O}(sT_{A})$ is the cost of forming the Gaussian sketch, $\mathcal{O}(ns\log s)$ is the cost of forming the SRTT sketch and $\mathcal{O}(s^{3})$ is the cost of computing the singular values of $\Gamma_{1}A\Gamma_{2}$.
@@ 2.4 Randomized norm estimation
In order to monitor how well our algorithm performs, we need an efficient way to estimate the norm of the error. Randomized norm estimation [26, 44] is an efficient method for finding an approximation to a norm of a matrix. In this work, we focus on the Frobenius norm as it is a natural choice for low-rank approximation and easier to estimate. Let $A\in\mathbb{R}^{m\times n}$ be a matrix and $\hat{A}_{r}\in\mathbb{R}^{m\times n}$ be a rank-$r$ approximation to $A$. Then we would like to estimate $\left\lVert A-\hat{A}_{r}\right\rVert_{F}$. As before, we can sketch to get an approximation to $A-\hat{A}_{r}$ and then compute the Frobenius norm. It turns out that the sketch size of $O(1)$, say $5$, suffices. More specifically, we have the following theorem from [26].
@@
This small sketch can be used to obtain an additional set of row and column indices by using pivoting on it, similarly to Algorithm 1. Specifically, we apply pivoting on the sketched residual $X-\hat{X}$ to get an extra set of $s$ column indices, denoted $J_{s}$. Next, we apply pivoting on the chosen columns of the residual matrix,
$A(:,J_{s})-A(:,J)A(I,J)^{\dagger}A(I,J_{s})$
to extract an extra set of $s$ row indices $I_{s}$. It is important to note that $I_{s}$ and $J_{s}$ will not include indices already chosen in $I$ and $J$ because the residual matrix $A-A(:,J)A(I,J)^{\dagger}A(I,:)$ is zero in the rows and columns corresponding to $I$ and $J$. The procedure for obtaining $s$ additional indices from the sketched residual is outlined in Algorithm 4.
Algorithm 4 Pivoting on the residual matrix
1:$X_{R}\in\mathbb{R}^{s\times n}$ sketch of the residual matrix, $I$ row indices, $J$ column indices
2:$I_{s}$ and $J_{s}$, extra sets of indices
3:function $[I_{s},J_{s}]=\mathtt{Pivot\_Residual}(X_{R},I,J)$
4:Apply CPQR on $X_{R}$. Let $J_{s}$ be the $s$ column pivots,
5:Set $C_{R}=A(:,J_{s})-A(:,J)A(I,J)^{\dagger}A(I,J_{s})\in\mathbb{R}^{m\times s}$
6:Apply CPQR on $C_{R}^{T}$. Let $I_{s}$ be the $s$ row pivots.
The complexity of Algorithm 4 is $\mathcal{O}((m+n)s^{2}+mrs+r^{3})=\mathcal{O}(n+mr+r^{3})$, where $r=\max\{|I|,|J|\}$ since $s=\mathcal{O}(1)$. The dominant cost lies in the computation of $A(:,J)A(I,J)^{\dagger}A(I,J_{s})$, which requires $\mathcal{O}(mr+r^{3})$ operations. By adding these additional indices, we can improve the accuracy of the CUR decomposition. Algorithm 4 will be applied later in AdaCUR to refine the approximation.
@@ 3 Proposed method
Let us first restate the problem. Let $A(t)\in\mathbb{R}^{m\times n}$ be a parameter-dependent matrix with $t\in D$ where $D\subseteq\mathbb{R}$ is a compact domain and $t_{1},t_{2},...,t_{q}\in D$ be a finite number of distinct sample points ordered in some way, e.g., $t_{1}<t_{2}<\cdots<t_{q}$ for $t\in\mathbb{R}$. Then find a low-rank approximation of $A(t_{1}),A(t_{2}),...,A(t_{q})$. [^7] Although our algorithms are presented for a finite number of sample points, they are likely to work on a continuum, e.g. by choosing the indices for the closest point $t_{i}$ for each $t\in[t_{1},t_{q}]$.
@@
The goal of this work is to devise an efficient algorithm that reuses the indices to their fullest extent in the CUR decomposition as we iterate through the parameter values. To see how we can use the same indices for different parameter values, recall (2) with oversampling from [50]:
(12) $\left\lVert A-A_{I\cup I_{0},J}\right\rVert_{F}\leq\left\lVert Q_{C}(I\cup I_{0},:)^{\dagger}\right\rVert_{2}\left\lVert V_{r}(J,:)^{-1}\right\rVert_{2}\left\lVert A-\llbracket{A}\rrbracket_{r}\right\rVert_{F}.$
@@
Here, $I_{0}$ is an extra set of row indices resulting from oversampling that are distinct from $I$.
The bound (12) leads us to two observations. First, as described earlier in the introduction, if $A(t)\in\mathcal{M}_{IJ}^{(r)}(\delta)$, then
$\left\lVert A-A_{I\cup I_{0},J}\right\rVert_{F}\leq\delta\left\lVert A-\llbracket{A}\rrbracket_{r}\right\rVert_{F}.$
Therefore, if $\delta$ is sufficiently small, say $\mathcal{O}(10)$, the sets of indices $I\cup I_{0}$ and $J$ yield a good CUR approximation. However, this does not imply that a high value of $\delta$ necessarily results in a poor CUR approximation. Let us illustrate this with a numerical example. We use the synthetic example from [9] given by
(13) $A(t)=e^{tW_{1}}e^{t}De^{tW_{2}},t\in[0,1]$
where $D\in\mathbb{R}^{n\times n}$ is diagonal with entries $d_{jj}=2^{-j}$ and $W_{1},W_{2}\in\mathbb{R}^{n\times n}$ are randomly generated skew-symmetric matrices. The singular values of $A(t)$ are $e^{t}2^{-j}$ for $j=1,2,...,n$. We take $n=500$ and take $101$ equispaced points in the interval $[0,1]$ for $t$. The matrix exponentials were computed using the $\mathtt{expm}$ command in MATLAB.
(a) Rel. error plot: CUR and trunc. SVD
(b) Gap between theory and practice of the CUR bound in (12).
Figure 3: Testing the gap between theory and practice for the CUR bound in (12) using the same set of indices for every parameter value. The parameter-dependent matrix in this example is the same as the one in Section 4.1 (Equation (13)). The target rank is $r=33$ in this experiment.
The results are shown in Figure 3. We begin with an initial set of indices and keep using the same set of indices for all other parameter values. The target rank is $33$ in this experiment. As shown in Figure 3a, despite using the same set of indices, we achieve a relative error of about $10^{-7}$ throughout, losing only about 11112222 digits of accuracy compared to the initial accuracy. On the other hand, in Figure 3b, we observe that the bound provided by (12) significantly overestimates the true bound. This demonstrates that the quantity $\left\lVert Q_{C}(I\cup I_{0},:)^{\dagger}\right\rVert_{2}\left\lVert V_{r}(J,:)^{-1}\right\rVert_{2}$ alone is insufficient to explain the effectiveness of the CUR decomposition. Such gaps between theoretical bounds and practical performance are common and have been observed in other works, such as [52].
In many problems, $\delta$ may be large, yet the CUR decomposition can provide a far better approximation than what is suggested by the bound in (12). Moreover, in most practical cases $\left\lVert A-\llbracket{A}\rrbracket_{r}\right\rVert_{F}$ is unknown. This observation motivates us to check whether the previously obtained indices can still be used before recalculating them entirely. To guard against potential large errors, we incorporate error and rank estimation in our algorithms, which we discuss in detail in Sections 3.2 and 3.3.
@@ 3.1 Computing indices from scratch
In the next two sections, we discuss our algorithms, AdaCUR and FastAdaCUR. For both algorithms, we need to start with an index set for the CUR decomposition. This can sometimes be done offline, for example, if we are solving a Matrix PDE, we can compute the initial set of indices for the CUR decomposition from the initial condition matrix. In this work, we use Algorithm 2 to get a set of indices from scratch. The procedure is to first approximate the $(\epsilon/\sqrt{n})$-rank of the initial matrix $A(t_{1})$ using randomized rank estimation and then apply pivoting on a random sketch with the estimated $(\epsilon/\sqrt{n})$-rank to get the initial set of indices. Here, we approximate the $(\epsilon/\sqrt{n})$-rank to ensure a relative error of $\epsilon$ for the CUR decomposition in the Frobenius norm. More specifically, the $1/\sqrt{n}$ factor arises from the worst case scenario where the trailing singular values of $A(t)$ are all equal. In such cases, the $(\epsilon/\sqrt{n})$-rank is required to guarantee a relative accuracy of $\epsilon$ in the Frobenius norm. If additional information about the spectrum of $A(t)$ is available–for instance, if $A(t)$ exhibits rapid singular value decay–then $\epsilon/\sqrt{n}$ can be replaced by $C\epsilon$, where $C<1$ is a constant. However, to guarantee accuracy, we use $(\epsilon/\sqrt{n})$-rank throughout work. The procedure for computing indices from scratch will be used for getting an initial set of indices for AdaCUR and FastAdaCUR, but also in AdaCUR when the error becomes too large that we need to recompute the indices altogether from scratch.
@@ 3.2 AdaCUR algorithm for accuracy
AdaCUR is an algorithm aimed at efficiently computing an accurate low-rank approximation of a parameter-dependent matrix with a given error tolerance. An overview of AdaCUR goes as follows. The algorithm starts by computing the initial set of indices for $A(t_{1})$ as discussed in Section 3.1. For subsequent parameter values, i.e., $A(t_{j})$ for $j=2,...,q$, we first verify whether the indices obtained from the previous parameter value are able to meet some given error tolerance, using an estimate of $\|A-CU^{\dagger}R\|_{F}$. If so, we continue to the next parameter value. Should the indices fail to meet the tolerance, we make low-cost minor modifications to the indices and test whether the adjustments satisfy the error tolerance. If this is still unsuccessful, we recompute the indices entirely from scratch and move onto the next parameter value. Minor modifications include adding a small fixed number of indices, reordering them based on their importance and removing some if necessary. The details are in the following paragraph. The addition and removal of indices, as well as recomputation of indices using rank estimation if necessary, make the algorithm rank-adaptive and computing the relative error at each instance make the algorithm certifiable. AdaCUR is presented in Algorithm 6.
@@
12:        $[I_{1},J_{1}]=\mathtt{Pivot\_Residual}(E_{s},I,J)$triangleright Algorithm 4
@@
14:        $[\sim,\sim,I_{2}]=\mathtt{sRRQR}\left(A(t_{j})\left(I\cup I_{0},J\right)^{T}\right)$ 888$[Q,R,I]=\mathtt{sRRQR}(A)$ performs strong rank-revealing QR factorization [27] on $A$ where $Q$ is an orthonormal matrix, $R$ is an upper triangular matrix and $I$ is a set of pivots.
@@
AdaCUR (Algorithm 6) takes the parameter-dependent matrix $A(t)$, evaluation points $t_{1},...,t_{q}\in D$, error tolerance $\epsilon$, error sample size $s$, and oversampling parameter $p$ as input. The algorithm produces the low-rank factors $C(t_{j}),R(t_{j}),U(t_{j})$ as output such that the factors are a subset of columns, rows and their intersection of $A(t_{j})$, respectively, and $A(t_{j})\approx C(t_{j})U(t_{j})^{\dagger}R(t_{j})$. If the entries of the parameter-dependent matrix can be computed quickly afterwards, the algorithm can be modified to output only row and column indices, saving storage. AdaCUR is quite involved, so we break it down into two core parts:
  • Section 3.2.1: Error estimation using previous indices (lines 55558888),
  • Section 3.2.2: Low-cost minor modifications (lines 1010101019191919).
The other lines involve computing the indices from scratch, which follow the same discussion as Section 3.1.
3.2.1 Initial error estimation